Two thirds of a number reduced by 11 is equal to 4 more than the number. Find the number.
Answer by checkley77(12844)

2X/3-11=X+4
2X/3-X=4+11
-X/3=15
-X=45
X=-45 ANS.
PROOF:
2*-45/3-11=-45+4
-90/3-11=-41
-30-11=-41
-41=-41